Armature of a 6-pole, 6-circuit D.C. shunt motor takes 400 A at a speed of 350 r.p.m. The flux per pole is 80 milli-webers, the number of armature turns is 600, and 3% of the


torque is lost in windage, friction and iron-loss. Calculate the brake-horse-power.



1
Expert's answer
2021-08-03T16:48:12-0400

Solution:-

Number of armature conducter=1200

Electromagnetic torque developed is

=T Nw-m

Armature power =Tw


"P=T\times2\pi\times\frac{350}{60}=36.67Twatt"

To calculate armature power in term of electrical parameters E must be known

"E=\frac{\phi zN}{60}\times\frac{P}{A}"


"E=80\times10^{-3}\times1200\times\frac{350}{60}\times\frac{6}{6}=560V"

With the current of 400 A armature power =500 "\times" 400 watt



"T=560\times\frac{400}{36.67}=6108..5Nw-m"

Since 3% of this torque is required for coming different loss term

Net torque "=0.97\times6180=5925Nw-m"

1hp=0.746kW

Where hp= horse power

Net output power in (kW) =

"5925\times36.67\times10^{-3}=217.27KW"

Converting in(brake horse power) BHP output =291.25HP